OpenStudy (anonymous):

how to represent (set of all complex numbers)^2? I know that set of all complex numbers is just a+bi

OpenStudy (anonymous):

thanks @Hoa

OpenStudy (anonymous):

a^2 +2abi -b^2 Is this what you're asking?

OpenStudy (turingtest):

the set of all complex numbers is usually represented by \(\mathbb C\)...

OpenStudy (anonymous):

somewhat, it is just a general form of (all set of complex number)^2\[C^2\]

OpenStudy (turingtest):

\(\mathbb C^2\) would be a 2-dimensional complex vector space

OpenStudy (anonymous):

Obtained by squaring (a+bi) (a+bi)(a+bi) = a^2 + abi +abi +b^2i^2 =a^2 +2abi +b^2*(-1) = a^2 + 2abi -b^2

OpenStudy (anonymous):

Yeah, what Turing said is the only other way to interpret this, from what I can tell. It would be the set of CxC basically, which means ordered pairs of complex numbers. The set would consist of (a+bi, c+di)

OpenStudy (sirm3d):

\[\large {a+bi\in \mathbb C\\(a+b)^2=(a^2-b^2)+(2ab)i\in \mathbb C}\]

OpenStudy (anonymous):

so who's right?

OpenStudy (sirm3d):

try posting the original question.

OpenStudy (anonymous):

What is the dimension of \[C^2\]as a real vector? @sirm3d

OpenStudy (sirm3d):

now, it's clear. @TuringTest is right.

OpenStudy (anonymous):

Unless you understand the question very well, paraphrasing what is asked is usually a bad idea. Very often, you'll word things in an unclear manner or leave out critical information.

OpenStudy (anonymous):

which is my weakness @SmoothMath